# SynCDR : Training Cross Domain Retrieval Models with Synthetic Data

### 31 Dec 2023 v1

@@ 4.3 Further Analyzing Synthetic Data
We find the least edit amount from Img2Img data, followed by InstructPix2Pix and CUT, while ELITE generates images that are the most distinct from source since it is not tied down to closely mimicking the source real image. These reflect the amount of edits we can see in examples in Fig. 2. In Appendix D, we present additional examples of synthetic data from each method. Additionally, factoring in results from Tab. 4, we find that larger edits corresponded to higher benefit of using PPP. In other words, the value of the pair label between a synthetic example and its real counterpart (indicating they are the same category) increases the more different the two are.
@@
5 Conclusion
Limitations. The best synthetic data for SynCDR uses textual prompts which describe the domains. This may not work if domains involved cannot be described this way. Here, a method like Textual Inversion [7] could be used to learn the properties of a domain as a new token embedding. We evaluated this method on DomainNet in Sec. A.1 but found its performance to be poorer than using textual descriptions. Experimenting on tasks where textually describing domains is not possible, is a topic of future work.
Appendix
Appendix A Additional Results
A.1 Using Textual Inversion to learn domains
In Sec. 5, under limitations, we mentioned that our synthetic data generation methods use simply a word or a phrase to represent the target domain (Tab. 12 lists these for different methods). This may not be possible in scenarios where this description is unavailable, e.g. in data collected from specific sensors. We hypothesize that in such a case, Textual Inversion [7] can be used. It allows us to encode a concept exemplified via a set of input images in a new token by optimizing its embeddings such that a pre-trained and frozen Stable Diffusion model generates the input images conditioned on this new token. Note that this description obfuscates some details to provide an intuitive understanding. For a full and precise description of textual inversion, we refer readers to Sec 3 of [7].
Method C-P C-S P-S Avg
SynCDR (ELITE) 45.4 ± 0.7 44.2 ± 0.4 46.4 ± 1.3 45.3 ± 0.9
Table 6: Using Textual Inversion to learn Domain Properties. We see that when we use a textually inverted token (see Sec. A.1 and [7]) that encodes domain properties for generating synthetic data with ELITE, cross-domain retrieval performance of SynCDR is similar or decreases. This happens possibly because Stable Diffusion’s understanding of common domains such as paintings, clipart and sketches is quite good. Additionally, as discussed in Sec. A.1, we see that in cases of generating paintings or clipart from sketches, using textual inversion is less accurate in preserving category of the input example. See examples in Fig. 3
From the results in Tab. 6, we see that textually inverting a domain’s properties leads to similar or poorer results overall. It is poorer in the case when sketch is involved as a domain. On digging deeper, we find that when generating paintings or cliparts from sketches, using textual inversion is poorer at preserving the category of the input sketch example. For instance, NCM accuracy (as defined in Sec. 4.3) of synthetic paintings generated with textual inversion is 0.54 compared to 0.61 without (see some examples of such cases in Fig. 3). This suggests the generation process somehow get skewed more to pay attention on only the style in such cases. We leave further investigation of this method to future work.
A.2 SynCDR performance with category overlap
@@
Table 7: Performance in the case of non-zero category overlap across domains in training data. In the introduction of the main paper, we motivated how synthetic data can come to the rescue in cases of missing similar category training examples across the two domains (i.e. 0% category overlap). Here, we see that synthetic data (used to train SynCDR) is still useful to improve the performance of CDS [18] even with non-zero category overlap.
@@
As discussed in the introduction, we can see the improved efficacy of CDS [18] as it gets training examples from same categories across the two domains. This is in constrast to ID which improved much less. We can see that SynCDR still helps improve over CDS performance significantly in this case.
A.3 Distilling large CLIP models for cross-domain retrieval
As mentioned in Sec. 4.1 and as shown in Tab. 8, image embeddings from large CLIP transformer models are robust across domains, but are expensive to obtain on smaller memory devices at inference time. In the top section of Tab. 8, we report the performances of two of these CLIP pre-trained models. The CLIP RN-50 is a larger version of the ResNet-50 model with $\sim$40M parameters compared to the latter’s $\sim$25M. The CLIP ViT-L is a large transformer model, which provides highly robust image embeddings at test time, but is expensive to store and run with $\sim$305M parameters.
@@ A.3 Distilling large CLIP models for cross-domain retrieval
where $D_{KL}$ is the KL divergence between two distributions. In essence, we attempt to distill the similarity predictions made by the teacher CLIP model $g$ for training images across domains (terms 1 and 3 of Sec. A.3), and those within the same domain (terms 2 and 4 of Sec. A.3).
@@
A.4 Full Retrieval Results
@@
Table 9: Cross Domain Retrieval performance (Prec@1, 5, 15) across different scenarios of DomainNet. Here we additionally report precisions at different number of retrieved examples for all models from Tab. 1.
@@
[…]
@@
CDS - 33.4 ± 0.7 28.1 ± 0.5 23.6 ± 0.7 45.9 ± 0.8 37.8 ± 0.6 32.2 ± 0.5 51.4 ± 0.8 44.4 ± 0.3 39.2 ± 0.3
ID - 36.0 ± 0.8 30.4 ± 0.4 25.6 ± 0.5 49.0 ± 0.8 40.8 ± 0.5 35.1 ± 0.3 54.9 ± 0.5 48.0 ± 0.3 42.5 ± 0.2
SynCDR (Ours) CUT 36.7 ± 1.0 31.4 ± 0.5 26.4 ± 0.4 48.8 ± 1.0 40.5 ± 0.4 34.9 ± 0.5 54.3 ± 1.4 48.0 ± 0.4 42.8 ± 0.5
Img2Img 36.3 ± 0.7 30.8 ± 0.8 26.3 ± 0.8 48.7 ± 0.8 41.5 ± 0.3 35.8 ± 0.3 54.5 ± 1.3 48.1 ± 0.9 43.1 ± 0.7
InstructPix2Pix 37.9 ± 0.9 32.6 ± 0.9 27.8 ± 0.9 49.6 ± 1.0 41.5 ± 0.4 35.8 ± 0.3 54.5 ± 1.1 48.4 ± 0.6 43.1 ± 0.6
ELITE 37.1 ± 1.0 32.6 ± 0.4 28.6 ± 0.2 49.1 ± 1.0 40.3 ± 0.7 34.8 ± 0.7 54.8 ± 1.1 48.8 ± 0.4 43.8 ± 0.4
@@
CDS - 38.9 ± 1.2 33.5 ± 0.9 27.0 ± 0.6 41.3 ± 0.7 35.3 ± 0.5 28.2 ± 0.2 63.3 ± 1.3 55.8 ± 0.9 45.2 ± 0.9
ID - 40.0 ± 1.0 35.3 ± 0.2 28.7 ± 0.3 43.5 ± 0.6 38.1 ± 0.5 30.2 ± 0.5 67.0 ± 0.6 59.9 ± 0.5 49.7 ± 0.3
SynCDR (Ours) CUT 41.1 ± 0.6 35.8 ± 0.7 29.0 ± 0.7 45.0 ± 0.8 39.7 ± 0.5 31.8 ± 0.2 67.5 ± 0.5 61.1 ± 0.3 51.3 ± 0.2
Img2Img 40.9 ± 0.8 35.8 ± 0.8 29.2 ± 0.7 43.5 ± 0.8 38.2 ± 0.6 30.2 ± 0.6 67.7 ± 0.4 61.4 ± 0.3 51.4 ± 0.4
InstructPix2Pix 42.8 ± 0.6 37.5 ± 0.8 30.8 ± 0.9 46.3 ± 1.1 40.5 ± 1.2 32.3 ± 1.4 67.5 ± 0.4 61.1 ± 0.3 51.2 ± 0.3
ELITE 43.6 ± 0.5 38.6 ± 0.6 31.6 ± 0.4 48.3 ± 0.4 41.9 ± 0.4 34.1 ± 0.3 67.5 ± 0.9 61.0 ± 0.7 51.3 ± 0.9
@@
CDS - 45.7 ± 1.0 39.1 ± 0.7 32.6 ± 0.6
ID - 48.4 ± 0.7 42.1 ± 0.4 35.3 ± 0.4
SynCDR (Ours) CUT 48.9 ± 0.9 42.7 ± 0.5 36.0 ± 0.5
Img2Img 48.6 ± 0.8 42.6 ± 0.7 36.0 ± 0.6
InstructPix2Pix 49.8 ± 0.9 43.6 ± 0.8 36.8 ± 0.8
ELITE 50.1 ± 0.9 43.9 ± 0.6 37.4 ± 0.5
Table 11: Cross Domain Retrieval performance (Prec@1, 5, 15) in different scenarios of Office-Home. Here we additionally report precisions at different number of retrieved examples for all models from Tab. 3.
@@
Appendix B Synthetic Data Generation Details
B.1 Contrastive Unpaired Translation (CUT)
@@ B.2 Img2Img/SDEdit
B.3 InstructPix2Pix
B.4 ELITE
Appendix C Other Training Details
For training SynCDR, we use most of the same hyperparameters as CDS [18]. Our model architecture is a ResNet-50 backbone followed by a fully connected layer resulting in an output feature dimension of 512. Before training, the backbone is initialized to Imagenet pre-trained weights. We use SGD with a learning rate of 0.003, with a batch size of 32 and momentum 0.9 for training SynCDR using a combination of the CDS and PPP losses as described in Sec. 3. We train for 15 epochs validating after each epoch and do early stopping based on validation set Prec@1.
Appendix D More Synthetic Generation Examples
Figs. 6, LABEL:, 7, LABEL: and 8 show some more examples of translations from the CUB and DomainNet datasets.

### 19 Mar 2024 v2

@@ 4.3 Further Analyzing Synthetic Data
We find the least edit amount from Img2Img data, followed by InstructPix2Pix and CUT, while ELITE generates images that are the most distinct from source since it is not tied down to closely mimicking the source real image. These reflect the amount of edits we can see in examples in Fig. 2. In Appendix 0.E, we present additional examples of synthetic data from each method. Additionally, factoring in results from Tab. 4, we find that larger edits corresponded to higher benefit of using PPP. In other words, the value of the pair label between a synthetic example and its real counterpart (indicating they are the same category) increases the more different the two are.
@@
4.4 SynCDR performance with category overlap
Table 6: Performance in the case of non-zero category overlap across domains in training data. In the introduction, we motivated how synthetic data can come to the rescue in cases of missing similar category training examples across the two domains (i.e. 0% category overlap). In this experiment performed using the DomainNet dataset, we see that synthetic data (used to train SynCDR) is still useful to improve the performance of CDS [22] even with non-zero category overlap.
@@ 4.4 SynCDR performance with category overlap
As discussed in the introduction, we can see the improved efficacy of CDS [22] as it gets training examples from same categories across the two domains. This is in contrast to in-domain ID which improved much less. We can see that SynCDR still helps improve over CDS performance significantly in this case.
5 Conclusion
Limitations. The best synthetic data for SynCDR uses textual prompts which describe the domains. This may not work if domains involved cannot be described this way. Here, a method like Textual Inversion [9] could be used to learn the properties of a domain as a new token embedding. We evaluated this method on DomainNet in Sec. 0.A.1 but found its performance to be poorer on average than using textual descriptions. Experimenting on tasks where textually describing domains is not possible, is a topic of future work.
Acknowledgments. This work was funded by the National Science Foundation and the Hariri Institute at Boston University.
Appendix 0.A Additional Results
0.A.1 Using Textual Inversion to learn domains
In the main paper under limitations, we mentioned that our synthetic data generation methods use simply a word or a phrase to represent the target domain (Tab. 12 lists these for different methods). This may not be possible in scenarios where this description is unavailable, e.g. in data collected from specific sensors. We hypothesize that in such a case, Textual Inversion [9] can be used. It allows us to encode a concept exemplified via a set of input images in a new token by optimizing its embeddings such that a pre-trained and frozen Stable Diffusion model generates the input images conditioned on this new token. Note that this description obfuscates some details to provide an intuitive understanding. For a full and precise description of textual inversion, we refer readers to Sec 3 of [9].
Table 7: Using Textual Inversion to learn Domain Properties. We see that when we use a textually inverted token (see Sec. 0.A.1 and [9]) that encodes domain properties for generating synthetic data with ELITE, cross-domain retrieval performance of SynCDR is similar or decreases. This happens possibly because Stable Diffusion’s understanding of common domains such as paintings, clipart and sketches is quite good. Additionally, as discussed in Sec. 0.A.1, we see that in cases of generating paintings or clipart from sketches, using textual inversion is less accurate in preserving category of the input example. See examples in Fig. 3
Method C-P C-S P-S Avg
SynCDR (ELITE) 45.4 ± 0.7 44.2 ± 0.4 46.4 ± 1.3 45.3 ± 0.9
From the results in Tab. 7, we see that textually inverting a domain’s properties leads to similar or poorer results overall. It is poorer in the case when sketch is involved as a domain. On digging deeper, we find that when generating paintings or cliparts from sketches, using textual inversion is poorer at preserving the category of the input sketch example. For instance, NCM accuracy (as defined in Sec. 4.3 of the main paper) of synthetic paintings generated with textual inversion is 0.54 compared to 0.61 without (see some examples of such cases in Fig. 3). This suggests the generation process somehow get skewed more to pay attention on only the style in such cases. We leave further investigation of this method to future work.
0.A.2 Distilling large CLIP models for cross-domain retrieval
As mentioned in Sec. 4.1 (main paper) and as shown in Tab. 8, image embeddings from large CLIP transformer models are robust across domains, but are expensive to obtain on smaller memory devices at inference time. In the top section of Tab. 8, we report the performances of two of these CLIP pre-trained models. The CLIP RN-50 is a larger version of the ResNet-50 model with $\sim$40M parameters compared to the latter’s $\sim$25M. The CLIP ViT-L is a large transformer model, which provides highly robust image embeddings at test time, but is expensive to store and run with $\sim$305M parameters.
@@ 0.A.2 Distilling large CLIP models for cross-domain retrieval
where $D_{KL}$ is the KL divergence between two distributions. In essence, we attempt to distill the similarity predictions made by the teacher CLIP model $g$ for training images across domains (terms 1 and 3 of Sec. 0.A.2), and those within the same domain (terms 2 and 4 of Sec. 0.A.2).
@@
0.A.3 Full Retrieval Results
Table 9: Cross Domain Retrieval performance (Prec@1, 5, 15) across different scenarios of DomainNet. Here we additionally report precisions at different number of retrieved examples for all models from Tab. 1 (main paper).
@@ 0.A.3 Full Retrieval Results
[…]
@@
Table 11: Cross Domain Retrieval performance (Prec@1, 5, 15) in different scenarios of Office-Home. Here we additionally report precisions at different number of retrieved examples for all models from Tab 3 (main paper).
@@
CDS - 36.5 ± 0.3 31.4 ± 0.2 26.7 ± 0.2 50.0 ± 0.4 42.1 ± 0.1 36.3 ± 0.2 54.9 ± 0.3 48.9 ± 0.1 43.6 ± 0.1
In-domain ID - 36.3 ± 0.3 31.2 ± 0.2 26.6 ± 0.1 50.0 ± 0.4 41.9 ± 0.1 36.1 ± 0.1 54.7 ± 0.3 48.9 ± 0.1 43.5 ± 0.1
SynCDR (Ours) CUT 38.3 ± 0.4 33.3 ± 0.2 28.7 ± 0.1 50.8 ± 0.2 42.6 ± 0.1 37.0 ± 0.1 55.5 ± 0.6 49.7 ± 0.2 44.7 ± 0.1
Img2Img 36.4 ± 0.3 31.4 ± 0.1 26.8 ± 0.1 49.9 ± 0.3 42.2 ± 0.2 36.4 ± 0.1 54.9 ± 0.5 49.2 ± 0.2 44.0 ± 0.2
InstructPix2Pix 38.6 ± 0.5 33.0 ± 0.2 28.6 ± 0.3 50.2 ± 0.2 42.3 ± 0.2 36.6 ± 0.2 54.7 ± 0.3 49.3 ± 0.1 43.9 ± 0.1
ELITE 38.3 ± 0.5 33.1 ± 0.2 29.1 ± 0.2 51.3 ± 0.3 42.5 ± 0.1 37.2 ± 0.1 55.4 ± 0.3 49.7 ± 0.2 44.8 ± 0.1
@@
CDS - 43.0 ± 0.2 38.0 ± 0.2 31.2 ± 0.1 45.2 ± 0.3 40.6 ± 0.2 32.5 ± 0.2 69.1 ± 0.2 62.5 ± 0.2 52.2 ± 0.2
In-domain ID - 43.4 ± 0.4 37.9 ± 0.2 31.0 ± 0.1 45.4 ± 0.4 40.4 ± 0.2 32.4 ± 0.2 69.0 ± 0.3 62.4 ± 0.1 52.1 ± 0.1
SynCDR (Ours) CUT 43.3 ± 0.5 38.0 ± 0.4 31.3 ± 0.3 46.6 ± 0.3 41.8 ± 0.3 34.5 ± 0.2 69.3 ± 0.3 62.8 ± 0.1 53.0 ± 0.1
Img2Img 43.4 ± 0.4 38.1 ± 0.2 31.2 ± 0.2 45.2 ± 0.4 40.4 ± 0.3 32.6 ± 0.3 69.0 ± 0.3 62.6 ± 0.1 52.5 ± 0.1
InstructPix2Pix 43.9 ± 0.2 39.2 ± 0.2 32.6 ± 0.2 46.7 ± 0.4 42.0 ± 0.2 34.0 ± 0.3 69.0 ± 0.2 62.5 ± 0.1 52.4 ± 0.1
ELITE 44.1 ± 0.6 39.3 ± 0.4 32.7 ± 0.4 46.3 ± 0.5 41.6 ± 0.3 33.8 ± 0.5 69.3 ± 0.3 62.9 ± 0.3 53.1 ± 0.4
@@
CDS - 49.8 ± 0.3 43.9 ± 0.2 37.1 ± 0.2
In-domainID - 49.8 ± 0.3 43.8 ± 0.2 36.9 ± 0.1
SynCDR (Ours) CUT 50.6 ± 0.4 44.7 ± 0.2 38.2 ± 0.2
Img2Img 49.8 ± 0.4 44.0 ± 0.2 37.3 ± 0.2
InstructPix2Pix 50.5 ± 0.3 44.7 ± 0.2 38.0 ± 0.2
ELITE 50.8 ± 0.4 44.8 ± 0.3 38.4 ± 0.3
@@
0.A.4 Feature Visualization
Fig. 4 shows t-SNE visualization of the output features from the network before and after SynCDR training. The two domains in the plot are Clipart and Painting from the DomainNet dataset and we used 1000 points from the test set of each domain picked at random for the visualization. We see that features get better clustered class-wise and are more aligned across domains after SynCDR training.
Figure 4: Feature visualization before and after training. The examples (from test set of DomainNet Clipart and Painting) get better clustered and more aligned across domains after training.
Appendix 0.B Synthetic Data Generation Details
0.B.1 Contrastive Unpaired Translation (CUT)
@@ 0.B.4 ELITE
Appendix 0.C Other Training Details
For training SynCDR, we use most of the same hyperparameters as CDS [22]. Our model architecture is a ResNet-50 backbone followed by a fully connected layer resulting in an output feature dimension of 512. Before training, the backbone is initialized to Imagenet pre-trained weights. We use SGD with a learning rate of 0.003, with a batch size of 32 and momentum 0.9 for training SynCDR using a combination of the CDS and PPP losses as described in Sec. 3 of the main paper. We trained all models for 15 epochs validating after each epoch and do early stopping based on validation set Prec@1.
Appendix 0.D Code
The code for running SynCDR has been attached in a zip folder with the supplement. We intend to make it publicly available on github with the final version of the paper.
@@ Appendix 0.D Code
Appendix 0.E More Synthetic Generation Examples
Figs. 7, LABEL:, 8, LABEL: and 9 show some more examples of translations from the CUB and DomainNet datasets.